A company invested $32,000, putting part of it into a savings account that earned 3.9% annual simple interest and the remainder in a stock fund that earned 12.9% annual simple interest. If the investments earned $1878 annually, how much was in each account?
savings account:
stock fund:

Answer by mananth(16946)

Part I 3.90% per annum ------------- Amount invested =x
Part II 12.90% per annum ------------ Amount invested = y
32000
Interest----- 1878.00

Part I 3.90% per annum ---x
Part II 12.90% per annum ---y
Total investment
x + 1 y= 32000 -------------1
Interest on both investments
3.90% x + 12.90% y= 1878
Multiply by 100
3.9 x + 12.9 y= 187800.00 --------2
Multiply (1) by -3.9
we get
-3.9 x -3.9 y= -124800.00
Add this to (2)
0 x 9 y= 63000
divide by 9
y = 7000
Part I 3.90% $ 25000
Part II 12.90% $ 7000

CHECK
25000 --------- 3.90% ------- 975.00
7000 ------------- 12.90% ------- 903.00
Total -------------------- 1878.00

m.ananth@hotmail.ca